OpenStudy (anonymous):

How to you find the measure of angle m<11 and m<12? m<11=4x m<12=2x-6

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

again, these two angles are supplementary

jimthompson5910 (jim_thompson5910):

so the measures of the angles add to 180

jimthompson5910 (jim_thompson5910):

m < 11 + m < 12 = 180 4x + (2x-6) = 180 4x + 2x - 6 = 180 Keep going to solve for x

OpenStudy (anonymous):

So 4x +2x-6 must equal 180.

OpenStudy (anonymous):

Then substitute x into 4x and 2x-6 to find the measures of the angles.

OpenStudy (anonymous):

m<11=124 m<12=56

jimthompson5910 (jim_thompson5910):

both are correct, nice job

OpenStudy (anonymous):

Thanks. So if both angle measures are supplemtary angles then I make the angle measures equal to 180?

jimthompson5910 (jim_thompson5910):

yes, if the angles are supplementary, then they add to 180

OpenStudy (anonymous):

Okay, I think I got it now.

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

What does a commplemtary angle look like?

jimthompson5910 (jim_thompson5910):

they add to 90 degrees

jimthompson5910 (jim_thompson5910):

so they form a right angle

OpenStudy (anonymous):

Ohh, okay. Thanks for all your help. You helped a lot.

jimthompson5910 (jim_thompson5910):

yw